OpenStudy (anonymous):

f(x)=x^5-2x+1......find the equation of the tangent line at x=1

OpenStudy (anonymous):

Is it 0?

OpenStudy (anonymous):

take the derivative, evaluate at 1 to find the slope, then use the point slope formula

OpenStudy (anonymous):

\[f(1)=0\] \[f'(x)=5x^4-2\] \[f'(1)=3\] use \[y-y_1=m(x-x_1)\] with \[x_1=1,y_1=0, m=3\]

OpenStudy (anonymous):

So i basically only did half of the equation?

OpenStudy (anonymous):

yeaaa but you still did a good job

OpenStudy (anonymous):

lol thnks, but i still feel like a fail...i wnt to finish the WHOLE equation. ;(